# ADAPT to Robustify Prompt Tuning Vision Transformers

### v1

@@ 5.2 False Sense of Security
In this section we evaluate the robustness of existing methods, combined with prompt tuning, against traditional and adaptive attacks. The results are presented in Tab. 1 and are consistent with what we discussed in Sec. 4.1. The results show that using traditional adversarial examples and adversarial training can lead to gradient obfuscation [1], and a false sense of security. Indeed, existing adversarial defense methods initially exhibit high performance when used with prompt tuning. However, as per [1], a sign of gradient obfuscation is that single step gradient based attacks are more effective than their multi-step counterparts. As expected, previous methods are more susceptible to single-step FGSM attacks than multi-step PGD attacks, and show significant vulnerability to our adaptive attack.
Params Method Clean ${}^{\dagger}$ ${}^{\dagger}$ $*{}^{*}$
@@
Table 1: Test set accuracy of previous adversarial robustness methods combined with prompt tuning against unperturbed inputs, traditional adversarial attacks, PGD10 (multi-step) and FGSM (single step) as well as our adaptive attack. All methods perform worse against FGSM than PGD10, a clear sign of gradient obfuscation. Additionally, all methods show very little robustness to the adaptive attack (Adaptive PGD10). ${}^{\dagger}$ indicates that the attack in question is non-adaptive while $*{}^{*}$ indicates an adaptive attack Eq. 13.
@@ 5.3 Adaptive Adversarial Training
We experiment with prompt tuning using previous SOTA adversarial defense methods, and compare to ADAPT. We present the results in Tab. 9. The results show that while previous methods show little to no robustness against adaptive adversarial attacks, ADAPT achieves competitive accuracy on unperturbed samples while showing considerable robustness to adaptive adversarial attacks. We provide additional results for different model sizes, ViT-L and ViT-S, in Tab. 3 and Tab. 4 respectively. The results show consistency with that of the ViT-B model.
For a fair comparison, we perform PT2 with the same number of tokens and tune the patch embedding as well as the linear classifier head for all methods, resulting in exactly the same number of tuned parameters. We include results of full model fine-tuning as an upper bound on performance, which requires the tuning of all the ViT parameters. We leave experimentation with PT and the effects of training the patch embedding to the ablation study in Tab. 8.
Method Params. Clean $*{}^{*}$ $*{}^{*}$ AA # of tuned params.
@@
ADAPT-CE (Ours) PT2+Emb 79.05 38.27 35.94 19.90 820K
ADAPT-KL (Ours) PT2+Emb 68.43 40.21 35.9 18.05 820K
@@
Table 2: Test set accuracy on the CIFAR10 dataset for different methods. The upper section corresponds to methods using prompt tuning, and the lower section corresponds to FT as an upper bound on performance. Existing methods show little to no robustness to adaptive attacks. In the prompt tuning scenario, our method shows substantial improvement in robustness compared to previous methods. $*{}^{*}$ indicates that the attack in question is adaptive Eq. 13.
Method Params. Clean PGD10 CW # of tuned parameters
@@
Table 3: CIFAR10 test set accuracy for the ViT/Large model.
Method Params. Clean PGD10 CW # of tuned parameters
TRADES PT2+Emb 86.06 3.98 3.81 740K
ADAPT+CE PT2 + Emb 73.7 34.33 31.52 740K
ADAPT+KL PT2 + Emb 63.37 36.1 31.74 740K
TRADES FT 81.1 51.71 45.16 57M
Table 4: CIFAR10 test set accuracy for the ViT/Small model
5.4 Beyond the Number of Parameters
We notice in our experiments that for adversarial training under the same conditions, PT2 significantly outperforms PT. We hypothesized that this is due to the fact that PT is simply augmenting the input to the frozen feature extractor $*\Phi*$ of the transformer which is vulnerable to adversarial examples, while PT2 inserts tokens in the intermediate layers of the feature extractor, modifying it directly. We perform an experiment to validate our intuition and confirm that the superior performance of PT2 over PT, is not simply due to an increased number of parameters (as with a fixed number of tokens per layer, PT2 requires more tuned parameters).
The results are presented in Tab. 5, and show a substantial increase in performance as we prompt more and more layers despite using the same number of tokens. This further validates our insight about needing to modify the feature extractor and shows that the performance improvement is not simply due to an increase in the number of tuned parameters.
Combination
Clean
PGD10
# of parameters
96$\times$1
27.33
18.34
81K
48$\times$2
32.74
21.07
81K
24$\times$4
40.03
24.83
81K
16$\times$6
40.78
25.49
81K
12$\times$8
41.14 25.68
81K
8$\times$12
40.24
25.17
81K
Table 5: Test set accuracy and the number of tuned parameters of different combinations of the same number of tokens. $m\times k$ means that we prompt the first $k$ layers directly with $m$ tokens each. Each setting is adopted during both training and testing. We measure against unperturbed samples and the adaptive PGD10 attack.
5.5 Imagenette and CIFAR100
We also evaluate our methods on Imagenette [13], which is a subset of ImageNet with 10 classes, and CIFAR100 [18], which consists of 100 classes of 32 $\times$ 32 images. Results are presented in Tab. 6 and Tab. 7 respectively, and show consistency with the CIFAR10 experiments.
Method Params Clean PGD10 CW # of tuned parameters
TRADES PT2+Emb 89.32 9.04 6.82 820K
ADAPT-CE PT2+Emb 59.95 31.34 28.25 820K
ADAPT-KL PT2+Emb 63.16 37.61 34.01 820K
TRADES FT 82.4 53.4 43.6 86M
Table 6: Test set accuracy on Imagenette
Method Params Clean PGD10 CW # of tuned parameters
TRADES PT2+Emb 78.84 5.54 3.52 820K
ADAPT-CE PT2+Emb 60.2 22.15 19.59 820K
ADAPT-KL PT2+Emb 50.63 21.77 17.56 820K
TRADES FT 65.5 32.89 28.93 86M
Table 7: Test set accuracy on CIFAR100
Method Params Clean PGD10 # of tuned params
ADAPT+CE PT 44.72 15.24 19K
ADAPT+KL PT 39.77 9.54 19K
ADAPT+CE PT2 67.45 26.75 230K
ADAPT+KL PT2 58.4 29.25 230K
ADAPT+CE PT+Emb 67.8 29.44 600K
ADAPT+KL PT+Emb 56.48 30.31 600K
ADAPT-CE PT2+Emb 79.05 38.27 820K
ADAPT-KL PT2+Emb 67.24 39.23 820K
6 Conclusion
7 Societal Impact
Supplementary Material: ADAPT to Robustify Prompt Tuning Vision Transformers
Masih Eskandar Tooba Imtiaz Zifeng Wang Jennifer Dy
8 Training Time Analysis
@@
With all that said, we provide training time details for the scenarios in Tab. 9 for those interested. We report the training time for each method in minutes. Each method was trained using an NVIDIA Tesla V100 SXM2.

### v2

@@ 5.2 False Sense of Security
In this section, we evaluate the robustness of existing methods, combined with prompt tuning, against traditional and adaptive attacks. The results are presented in table 1 and are consistent with what we discussed in adaptive attacks section. The results show that using traditional adversarial examples and adversarial training can lead to gradient obfuscation (Athalye et al., 2018), and a false sense of security. Indeed, existing adversarial defense methods initially exhibit high performance when used with prompt tuning. However, as per (Athalye et al., 2018), a sign of gradient obfuscation is that single-step gradient-based attacks are more effective than their multi-step counterparts. As expected, previous methods are more susceptible to single-step FGSM attacks than multi-step PGD attacks and show significant vulnerability to our adaptive attack.
Table 1: Test set accuracy of previous adversarial robustness methods combined with prompt tuning against unperturbed inputs, traditional adversarial attacks, PGD10 (multi-step) and FGSM (single step) as well as our adaptive attack. All methods perform worse against FGSM than PGD10, a clear sign of gradient obfuscation. Additionally, all methods show very little robustness to the adaptive attack (Adaptive PGD10). indicates that the attack in question is non-adaptive while indicates an adaptive attack eq. 13.
Params Method Clean PGD FGSM PGD
@@ 5.3 Adaptive Adversarial Training
We experiment with prompt tuning using previous SOTA adversarial defense methods and compare them to ADAPT. We present the results in table 2. The results show that while previous methods show little to no robustness against adaptive adversarial attacks, ADAPT achieves competitive accuracy on unperturbed samples while showing considerable robustness to adaptive adversarial attacks. We provide additional results for different model sizes, ViT-L and ViT-S, in the appendix, the results are consistent with that for ViT-B.
We also evaluate our methods on Imagenette (Howard, ), which is a subset of ImageNet with 10 classes, and CIFAR100 (Krizhevsky et al., b), which consists of 100 classes of 32 $\times$ 32 images. Results are presented in table 3 and table 4 respectively, and show consistency with the CIFAR10 experiments.
Note that for a fair comparison, we perform PT2 with the same number of tokens and tune the patch embedding and the linear classifier head for all methods, resulting in exactly the same number of tuned parameters. We include the results of full model fine-tuning (FT) as an upper bound on performance, which requires tuning all the ViT parameters. We conduct experimentation with PT and the effects of training the patch embedding in the ablation study (table 7). The results confirm the effectiveness of our design choices. Furthermore, we investigate the impact of the number of tuned parameters in the following sections.
Table 2: Test set accuracy on the CIFAR10 dataset for different methods. The upper section corresponds to methods using prompt tuning, and the lower section corresponds to FT as an upper bound on performance. Existing methods show little to no robustness to adaptive attacks. In the prompt tuning scenario, our method shows substantial improvement in robustness compared to previous methods. indicates that the attack in question is adaptive (eq. 13).
Method Params. Clean PGD10 CW AA # params
@@
MART PT2+Emb 52.28 12.1 9.62 15.25 820K
NFGSM PT2+Emb 77.95 0.1 0.6 2.03 820K
$\text{ADAPT}_{CE}$ PT2+Emb 79.05 38.27 35.94 19.90 820K
$\text{ADAPT}_{KL}$ PT2+Emb 68.43 40.21 35.9 18.05 820K
@@
Table 3: Test set accuracy on Imagenette
Method Params Clean PGD10 CW AA # params
TRADES PT2+Emb 89.32 9.04 6.82 7.08 820K
MART PT2+Emb 44.2 17.35 14.03 19.36 820K
NFGSM PT2+Emb 71.82 2.08 2.08 5.68 820K
$\text{ADAPT}_{CE}$ PT2+Emb 59.95 31.34 28.25 13.81 820K
$\text{ADAPT}_{KL}$ PT2+Emb 63.16 37.61 34.01 10.96 820K
TRADES FT 82.4 53.4 43.6 16.33 86M
Table 4: Test set accuracy on CIFAR100
Method Params Clean PGD10 CW AA # params
TRADES PT2+Emb 78.84 5.54 3.52 3.4 820K
MART PT2+Emb 29.29 6.59 4.99 8.5 820K
NFGSM PT2+Emb 57.91 0.6 0.9 1.67 820K
$\text{ADAPT}_{CE}$ PT2+Emb 60.2 22.15 19.59 7.92 820K
$\text{ADAPT}_{KL}$ PT2+Emb 50.63 21.77 17.56 5.94 820K
TRADES FT 65.5 32.89 28.93 9.99 86M
5.4 Evaluation against Blackbox Attacks
In this section, we evaluate the robustness of ADAPT and existing methods against black-box attacks in the prompt-tuning paradigm. We evaluate against two SOTA black-box attacks: one query-based attack, RayS (Chen & Gu, 2020) and one transfer-based ensemble attack, AdaEA (Chen et al., 2023b). Table 5 shows that ADAPT is robust to both the query-based and transfer-based attacks, while AT and TRADES (with prompt tuning) only show robustness to the transfer-based attack and are not robust to the query-based attack. We hypothesize that the reason is AT and TRADES (with prompt tuning) are trained on attacks generated using the gradient of the naturally trained frozen ViT. The transfer-based attack also relies on the gradients of naturally pre-trained models to construct its attack, while the query-based attack eliminates the need to use gradients. This aligns with our discussion of gradient obfuscation. Furthermore, against existing defenses, both black-box attacks are significantly more successful than white-box non-adaptive attacks (see table 2), which is another sign of gradient obfuscation in existing defenses when using prompt tuning.
Table 5: Test-set accuracy against black-box attacks, same training settings as Table 2.
Method Params RayS AdaEA
AT P2T+Emb 5.06 66.55
TRADES P2T+Emb 11.71 66.27
$\text{ADAPT}_{CE}$ P2T+Emb 43.05 58.21
$\text{ADAPT}_{KL}$ P2T+Emb 40.09 52.06
TRADES FT 56.78 61.73
5.5 Beyond the Number of Parameters
We notice in our experiments that for adversarial training under the same conditions, PT2 significantly outperforms PT. We hypothesized that this is due to the fact that PT is simply augmenting the input to the frozen feature extractor $\Phi*$ of the transformer which is vulnerable to adversarial examples, while PT2 inserts tokens in the intermediate layers of the feature extractor, modifying it directly. We perform an experiment to validate our intuition and confirm that the superior performance of PT2 over PT, is not simply due to an increased number of parameters (as with a fixed number of tokens per layer, PT2 requires more tuned parameters).
The results are presented in table 6, and show a substantial increase in performance as we prompt more and more layers despite using the same number of tokens. This further validates our insight about needing to modify the feature extractor and shows that the performance improvement is not simply due to an increase in the number of tuned parameters.
Table 6: Test set accuracy and the number of tuned parameters of different combinations. $m\times k$ means we prompt the first $k$ layers directly with $m$ tokens each. Each setting is adopted during both training and testing. We measure against unperturbed samples and the adaptive PGD10 attack.
Combination Clean PGD10 # params
96$\times$1 27.33 18.34 81K
48$\times$2 32.74 21.07 81K
24$\times$4 40.03 24.83 81K
16$\times$6 40.78 25.49 81K
12$\times$8 41.14 25.68 81K
8$\times$12 40.24 25.17 81K
Method Params Clean PGD10 # params
$\text{ADAPT}_{CE}$ PT 44.72 15.24 19K
$\text{ADAPT}_{KL}$ PT 39.77 9.54 19K
$\text{ADAPT}_{CE}$ PT2 67.45 26.75 230K
$\text{ADAPT}_{KL}$ PT2 58.4 29.25 230K
$\text{ADAPT}_{CE}$ PT+Emb 67.8 29.44 600K
$\text{ADAPT}_{KL}$ PT+Emb 56.48 30.31 600K
$\text{ADAPT}_{CE}$ PT2+Emb 79.05 38.27 820K
$\text{ADAPT}_{KL}$ PT2+Emb 67.24 39.23 820K
6 Conclusion
7 Broader Impact Statement
8 Acknowledgments
This project was supported by NIH grants R01CA240771 and U24CA264369 from NCI. We would like to thank the TMLR editors and reviewers as well as our colleagues and friends for helping improve this work.
Appendix A Different Model Sizes
We provide results for the experiment setting of table 1, for different ViT configurations, namely ViT-S in table 8 and ViT-L in table 9. The results are consistent with that of table 1.
Table 8: CIFAR10 test set accuracy for the ViT/Small model
Method Params. Clean PGD CW # params
TRADES PT2+Emb 86.06 3.98 3.81 740K
ADAPT+CE PT2 + Emb 73.7 34.33 31.52 740K
ADAPT+KL PT2 + Emb 63.37 36.1 31.74 740K
TRADES FT 81.1 51.71 45.16 57M
Table 9: CIFAR10 test set accuracy for the ViT/Large model.
Method Params. Clean PGD CW # params
@@ Appendix A Different Model Sizes
Appendix B Training Time Analysis
@@
With all that said, we provide training time details for the scenarios in table 10 for those interested. We report the training time for each method in minutes. Each method was trained using an NVIDIA Tesla V100 SXM2.